a, b please. thanks.
12) Given a) w = two vectors:u=(2,5), v= <3,-4%. Find Proj v b) write u as the sum of W, and Wz


12) Given two vectors:u=(2,5), v=<3,-4%. Find a) wi= Proj v b) write u as the sum of Wi and Wz

just A if possible please
Answer #1

c ε 10 ay 0 = (215), v=<3,-4) llyll=532+ (-4)2 = √9+16 = 125 -5. 2. W, = Pooj u - u yon V = (243) 45*]<37-4) = (6=20]<3,-4)